OpenStudy (anonymous):

What is the exact value of the square root of the quantity 128/14.? Simplify if possible.

OpenStudy (anonymous):

\[\frac{ \sqrt{128} }{ \sqrt{14} }\]

OpenStudy (anonymous):

answers are: \[\frac{ \8\sqrt{2} }{ \sqrt{2} }]\] \[\frac{ 8\sqrt{7} }{ \sqrt{7} }\]

OpenStudy (anonymous):

\[\frac{ 16\sqrt{2} }{ \sqrt{2} }\] \[\frac{ 16\sqrt{7} }{\sqrt{7} }\]

OpenStudy (anonymous):

my answer was the first but im not sure